OpenStudy (anonymous):

what is the solution to the equation x - 12 = 3 A. X = -9 B. X = 9 C. X = 15

OpenStudy (anonymous):

it should be a positive 15

OpenStudy (anonymous):

Oh yes positive sorry

OpenStudy (anonymous):

then it is c

OpenStudy (pawanyadav):

Add 12 both side

OpenStudy (zpupster):

x - 12 = 3 add 12 to both sides x-12+12=12+3 this how you do these types of problems